OpenStudy (anonymous):

x^2/x^2+3 * x+10/x multiply and simplify to lowest terms

OpenStudy (anonymous):

Did you mean \[\frac{ x ^{2} }{ x ^{2}+3} \times \frac{ x+10 }{ x }\]

OpenStudy (anonymous):

multiplying the top you get what? remember to times x^2 by both x and 10

OpenStudy (anonymous):

@Sky_Pirate

OpenStudy (anonymous):

... im not going to give you the answer.

OpenStudy (anonymous):

i was not asking yu for it love

OpenStudy (anonymous):

I mean, this is basic algebra. x^2 times by x is x^3 because you add the exponents. what are you having trouble with?

OpenStudy (anonymous):

i want others to work together and find it

OpenStudy (anonymous):

just an example

OpenStudy (anonymous):

uhh what.

OpenStudy (anonymous):

\[\frac{ x(x+10) }{ x ^{2}+3 }\] xD

OpenStudy (anonymous):

and how did you get that answer?

OpenStudy (anonymous):

you know ^^ first you factor top and the bottom you put together since that can't factor